OpenStudy (ray1998):

Find all the zeros of the equation. -3x^4 + 27x^2 + 1200 = 0 Fan and medal will be given to whomever can help! Thank you:)

OpenStudy (curiousquestions):

All of the zeros?

OpenStudy (curiousquestions):

What do you mean?

OpenStudy (ray1998):

Idk, that's what the problem is asking.

OpenStudy (tkhunny):

Can you find all the zeros of this? -3x^2 + 27x + 1200 = 0

OpenStudy (ray1998):

No, I'm behind in my class and don't know what to do. I'm trying my hardest just to catch up.

OpenStudy (curiousquestions):

Ok. Um...

OpenStudy (tkhunny):

First, and maybe it's not obvious, is everything is divisible by 3 or -3. -3x^2 + 27x + 1200 = 0 x^2 - 9x - 400 = 0 Can you solve this quadratic equation? Completing the Square? Factoring? Quadratic Equation? Graphing?